22)
solve: tan3x(tanx-1)=0

Answer by stanbon(57979)

solve: tan3x(tanx-1)=0
tan(3x) = 0 or tan(x)-1 = 0
3x = 0 or tan(x) = 1
x = 0 or x = pi/4 or x = (5/4)pi
====================
Cheers,
Stan H.